OpenStudy (anonymous):

Pre-calculus help! Limits of functions 1. Estimate the limit. ((sqrt x+2)-(sqrt 2)) / x. 2. Estimate the limit. ((sqrt 1-x) -1)/x

OpenStudy (loser66):

\[\lim_{x \rightarrow ?}\frac{\sqrt{x}+2-\sqrt{x}}{x}\]

OpenStudy (loser66):

x goes to where? 2 is under the square root or not?

OpenStudy (anonymous):

Yes, 2 is under the sqare root! * x goes to 0. Under the 2nd square root on the top is a 2 though.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ \sqrt{x+2} - \sqrt{2}}{ x }\]

OpenStudy (anonymous):

try multiplying by the numerator's conjugate so to speak